# Disentangling transients and their host galaxies with scarlet2: A framework to forward model multi-epoch imaging

### v1

@@ Abstract
Many science cases for wide-field time-domain surveys rely on accurate identification and characterization of the galaxies hosting transient and variable objects. In the era of the Legacy Survey of Space and Time (LSST) at the Vera C. Rubin Observatory the number of known transient and variable sources will grow by orders of magnitude, and many of these sources will be blended with their host galaxies and neighboring galaxies. A diverse range of applications – including the classification of nuclear and non-nuclear sources, identification of potential host galaxies in deep fields, extraction of host galaxy SEDs without requiring a transient-free reference image, and combined analysis of photometry from multiple surveys – will benefit from a flexible framework to model time-domain imaging of transients. We describe a time-domain extension of the scarlet2 scene modeling code for multi-epoch, multi-band, and multi-resolution imaging data to extract simultaneous transient and host galaxy models. scarlet2 leverages the benefits of data-driven priors on galaxy morphology, is fully GPU compatible, and can jointly model multi-resolution data from ground and space-based surveys. We demonstrate the method on simulated LSST-like supernova imaging, low-resolution Zwicky Transient Facility imaging of tidal disruption events, and Hyper Suprime Cam imaging of variable AGN out to $z=4$ in the COSMOS fields. We show that scarlet2 models provide accurate transient and host galaxy models as well as accurate measurement of host–transient spatial offsets, and demonstrate future applications to the search for ‘wandering’ massive black holes.
@@
journal: Astronomy &\&& Computing
@@ 1 Introduction
Time-domain surveys frequently employ methods based on a technique called ‘difference imaging’ to detect transient phenomena. In such an approach, a deep ‘reference image’ is produced from a stack of high quality images and subtracted from new images to produce a difference image (Alard and Lupton, 1998). Difference imaging enables the efficient detection and optimal extraction of multi-epoch fluxes of transient sources. In this work, we are primarily concerned with enabling follow-up investigations of the transient and its host once the transient has been identified by difference imaging. We propose a model of a variable point source and a non-varying host galaxy that is rendered to match all observations. Our approach is particularly advantageous when the difference imaging introduces ambiguities – for example, when the variable object is already contributing to the reference image, which is always the case for AGNs – or when simultaneously modeling the transient and the host is helpful, such as when measuring small spatial offsets between the host center and the transient. Our approach can also make use of imaging data from different surveys, enabling the automatic production of combined light curves without concerns about mismatches between the reference image epochs.
@@
Multi-resolution scene modeling methods have previously been applied to produce galaxy catalogs from deep stacks for the DESI Legacy Imaging Surveys (Dey et al., 2019), and to produce deep forced photometry of 400 million sources in low resolution imaging from the Wide-Field Infrared Survey Explorer, exploiting higher resolution SDSS imaging to anchor the photometry (Lang et al., 2016). However, scene modeling has rarely been implemented for multi-epoch imaging of transients. A version of the scene modeling approach was applied to produce supernova photometry in the Sloan Digital Sky Survey-II Supernova survey (Holtzman et al., 2008) and the Dark Energy Survey Supernova Program survey (Brout et al., 2019), but is yet to be applied at scale for nuclear transients and AGN.
@@
Scene modeling is particularly powerful when imaging data from ground and space-based telescopes is combined so that the high-resolution imaging with small PSF sizes can improve galaxy deblending solutions for sources that are marginally resolved in low-resolution ground-based imaging (Melchior et al., 2021). Software packages that can simultaneously model multi-band, multi-resolution imaging data include The Tractor (Lang et al., 2016), scarlet (Melchior et al., 2018), and Astrophot (Stone et al., 2023), the latter of which is GPU-accelerated. In contrast to The Tractor and Astrophot, scarlet is a non-parametric method: it does not require the adoption of an analytical profile such as a Sérsic profile for extended sources. Simple requirements on galaxy shape, such as a monotonically decreasing or symmetric profile, or positive flux in all regions of the galaxy, are enforced by proximal operators. Gradient descent with Autograd is used to fit the source parameters, with the differing colors of overlapping sources being the most important distinguishing factor for deblending.
@@
In this paper, we introduce a new time-domain extension to scarlet2 that enables modeling of multi-band, multi-resolution and multi-epoch imaging data, where we treat some sources to as variable and others as static. We test the methodology on simulated multi-epoch, multi-band supernova imaging, and then demonstrate it on real survey data for samples of a) TDEs and b) AGNs. For our two case studies, we apply our scene-modeling technique to a) high cadence, low resolution, shallow multi-epoch imaging from the Zwicky Transient Facility (ZTF; Graham et al., 2019; Bellm et al., 2019; Dekany et al., 2020) and b) low cadence, high resolution, deep multi-epoch imaging from the Hyper Suprime Cam Subaru Strategic Program (HSC-SSP) Transient Survey (Yasuda et al., 2019). We demonstrate the application of non-parametric galaxy morphology models, guided by data-driven neural network priors, to produce full scene models with accurate transient photometry and position measurements.
@@ 4.1 Background and motivation
Figure 5: Comparison of light curves of example TDE ZTF19abhhjc produced by forced photometry on ZTF difference imaging (left) and scarlet2 scene modeling (right).
@@ 4.2 Light curves, host galaxy morphology models, and spatial offsets for ZTF TDEs
In Figure 5 we show the the single-epoch fluxes measured by scarlet2 for one of the ZTF TDEs, ZTF19abhhjc, as well as the equivalent light curve produced using the forced photometry pipeline for ZTF difference imaging (Masci et al., 2019). scarlet2 is able to produce a light curve consistent with the difference imaging pipeline, and correctly assigns flux to the host galaxy and the nuclear transient where they are blended. We specifically chose ZTF19abhhjc because it represents a more difficult modeling problem: the host galaxy is compact (Figure 4), such that there is little extended emission beyond the nucleus to assist in decomposing the nuclear transient and host galaxy SEDs.
@@ 5.1 Background and motivation
As part of our HSC AGN analysis, we undertake a pilot study for the search for rare MBHs that are spatially offset from their host galaxy nuclei, in preparation for a larger-scale analysis with LSST. Recent cosmological simulations that do not artificially tie SMBHs to halo centers have found that the majority of SMBHs of mass $M_{\text{BH}}<10^{7}M_{\odot}$ have a $>15$kpc spatial offset from their host nucleus, with the fraction increasing for decreasing mass (Bellovary et al., 2019; Ricarte et al., 2021). Variability-selection has proven to be particularly effective at identifying low mass AGN in dwarf galaxies where traditional spectroscopic diagnostics are less effective (e.g. Baldassare et al., 2018, 2020; Ward et al., 2022; Burke et al., 2022), making it a promising approach for the identification of wandering MBHs. Forecasts for the number of $10^{4}M_{\odot}<M_{\text{BH}}<10^{6}M_{\odot}$ MBHs with detectable variability in LSST range from 1,500 to 21,000 depending on seed mass and wandering fraction (Burke et al., 2023).
There are a variety of predictions for the accretion luminosities of such wandering AGN (Pacucci et al., 2018; Di Matteo et al., 2023), but even if they are not observable from typical gas accretion, they may be observable as tidal disruption events if the wandering SMBHs retain their nuclear star clusters (e.g Lin et al., 2018). The goal of this pilot study is to determine what spatial offsets will be detectable by applying multi-epoch scene modeling to LSST-like imaging data, using a sample of variable AGNs with well-determined spectroscopic redshifts and black hole masses (Aihara et al., 2022; Burke et al., 2024).
@@ 5.2 Light curves, host galaxy SEDs, and spatial offset distributions for variable AGN in HSC
In Figure 7, we show a HSC observation, the scarlet2 model for that observation derived from fitting the full multi-epoch data set, and the corresponding residual image for two COSMOS AGN: one in a compact, AGN-dominated galaxy – which is very typical for this sample – and one in an extended galaxy of comparable flux to the variable AGN. scarlet2 produces high quality models and light curves for all AGN in the HSC sample. Three example light curves from this procedure are shown in Figure 8.
@@ 6 Summary and conclusions
This method will be most powerful when high cadence time-domain ground-based imaging from surveys like ZTF and LSST is jointly modeled with higher resolution space-based imaging from HST, Euclid (Racca et al., 2016), and the Roman Space Telescope (Akeson et al., 2019). The option for GPU acceleration means that image processing with scarlet2 can be run at scale. We have made the scarlet2 code available for the wider astrophysical community at https://github.com/pmelchior/scarlet2. Example jupyter notebooks for the methods described in each section are available at https://github.com/charlotteaward/TimeDomainScarlet2.

### v2

@@ Abstract
Many science cases for wide-field time-domain surveys rely on accurate identification and characterization of the galaxies hosting transient and variable objects. In the era of the Legacy Survey of Space and Time (LSST) at the Vera C. Rubin Observatory the number of known transient and variable sources will grow by orders of magnitude, and many of these sources will be blended with their host galaxies and neighboring galaxies. A diverse range of applications – including the classification of nuclear and non-nuclear sources, identification of potential host galaxies in deep fields, extraction of host galaxy spectral energy distributions without requiring a transient-free reference image, and combined analysis of photometry from multiple surveys – will benefit from a flexible framework to model time-domain imaging of transients. We describe a time-domain extension of the scarlet2 scene modeling code for multi-epoch, multi-band, and multi-resolution imaging data to extract simultaneous transient and host galaxy models. scarlet2 leverages the benefits of data-driven priors on galaxy morphology, is fully GPU compatible, and can jointly model multi-resolution data from ground and space-based surveys. We demonstrate the method on simulated LSST-like supernova imaging, low-resolution Zwicky Transient Facility imaging of tidal disruption events, and Hyper Suprime Cam imaging of variable AGN out to $z=4$ in the COSMOS fields. We show that scarlet2 models provide accurate transient and host galaxy models as well as accurate measurement of host–transient spatial offsets, and demonstrate future applications to the search for ‘wandering’ massive black holes.
@@
journal: Astronomy &\&& Computing (accepted)
@@ 1 Introduction
Time-domain surveys frequently employ methods based on a technique called ‘difference imaging’ to detect transient phenomena. In such an approach, a deep ‘reference image’ is produced from a stack of high quality images and subtracted from new images to produce a difference image (Alard and Lupton, 1998). Difference imaging enables the efficient detection and optimal extraction of multi-epoch fluxes of transient sources. In this work, we are primarily concerned with enabling follow-up investigations of the transient and its host once the transient has been identified by difference imaging. We propose a model of a variable point source and a non-varying host galaxy that is rendered to match all observations. We refer to this as a scene modeling approach, whereby all sources in the scene are modeled in a framework that incorporates information about observation parameters such as the world coordinate system, the point spread function, and the background noise. Our approach is particularly advantageous when the difference imaging introduces ambiguities – for example, when the variable object is already contributing to the reference image, which is always the case for active galaxy nuclei (AGNs) – or when simultaneously modeling the transient and the host is helpful, such as when measuring small spatial offsets between the host center and the transient. Our approach can also make use of imaging data from different surveys, enabling the automatic production of combined light curves without concerns about mismatches between the reference image epochs.
@@
Multi-resolution scene modeling methods have previously been applied to produce galaxy catalogs from deep stacks for the DESI Legacy Imaging Surveys (Dey et al., 2019), and to produce deep forced photometry of 400 million sources in low resolution imaging from the Wide-Field Infrared Survey Explorer, exploiting higher resolution SDSS imaging to anchor the photometry (Lang et al., 2016). However, scene modeling has rarely been implemented for multi-epoch imaging of transients. A version of the scene modeling approach was applied to produce supernova photometry in the Sloan Digital Sky Survey-II Supernova survey (Holtzman et al., 2008) and the Dark Energy Survey Supernova Program survey (Brout et al., 2019). The method is yet to be applied for nuclear transients and AGN, where degeneracies between emission from the galaxy nucleus and the AGN make deblending particularly challenging. Current best practice is to obtain a single epoch of high-resolution space-based imaging, and then to fit a galaxy profile and a nuclear point source to that image (e.g. Kim2008; Zhong et al., 2022). In such cases, a prior on the AGN SED can be obtained by fitting AGN and galaxy templates to a spectrum of the galaxy nucleus. Convolutional neural networks have also been used for AGN-host decomposition of James Webb Space Telescope imaging (Margalef2024). To our knowledge, forward modeling of variable AGN across multi-epoch ground-based imaging has yet to be used as a method for AGN-host decomposition or extraction of AGN light curves.
@@
Scene modeling is particularly powerful when imaging data from ground and space-based telescopes is combined so that the high-resolution imaging with small PSF sizes can improve galaxy deblending solutions for sources that are marginally resolved in low-resolution ground-based imaging (Melchior et al., 2021). Software packages that can simultaneously model multi-band, multi-resolution imaging data include The Tractor (Lang et al., 2016), scarlet (Melchior et al., 2018), and Astrophot (Stone et al., 2023), the latter of which is GPU-accelerated. In contrast to The Tractor and Astrophot, scarlet is a non-parametric method: it does not require the adoption of an analytical profile such as a Sérsic profile for extended sources. Simple requirements on galaxy shape, such as a monotonically decreasing or symmetric profile, or positive flux in all regions of the galaxy, are enforced by proximal operators. Gradient descent with Autograd is used to fit the source parameters, with the differing colors of overlapping sources being the most important distinguishing factor for deblending. The assumption that the morphology is constant across bands provides a strong constraint that is essential when modeling crowded galaxy fields. For large galaxies with color gradients, multi-component galaxy models enable accurate deblending while working under this constraint.
@@
In this paper, we introduce a new time-domain extension to scarlet2 that enables modeling of multi-band, multi-resolution and multi-epoch imaging data, where we treat some sources as variable and others as static. We test the methodology on simulated multi-epoch, multi-band supernova imaging, and then demonstrate it on real survey data for samples of a) TDEs and b) AGNs. For our two case studies, we apply our scene-modeling technique to a) high cadence, low resolution, shallow multi-epoch imaging from the Zwicky Transient Facility (ZTF; Graham et al., 2019; Bellm et al., 2019; Dekany et al., 2020) and b) low cadence, high resolution, deep multi-epoch imaging from the Hyper Suprime Cam Subaru Strategic Program (HSC-SSP) Transient Survey (Yasuda et al., 2019). We demonstrate the application of non-parametric galaxy morphology models, guided by data-driven neural network priors, to produce full scene models with accurate transient photometry and position measurements.
@@ 4.1 Background and motivation
Figure 5: Comparison of light curves of example TDE ZTF19abhhjcc produced by forced photometry on ZTF difference imaging (left) and scarlet2 scene modeling (right).
@@ 4.2 Light curves, host galaxy morphology models, and spatial offsets for ZTF TDEs
In Figure 5 we show the the single-epoch fluxes measured by scarlet2 for one of the ZTF TDEs, ZTF19abhhjcc, as well as the equivalent light curve produced using the forced photometry pipeline for ZTF difference imaging (Masci et al., 2019). scarlet2 is able to produce a light curve consistent with the difference imaging pipeline, and correctly assigns flux to the host galaxy and the nuclear transient where they are blended. We specifically chose ZTF19abhhjcc because it represents a more difficult modeling problem: the host galaxy is compact (Figure 4), such that there is little extended emission beyond the nucleus to assist in decomposing the nuclear transient and host galaxy SEDs.
@@ 5.1 Background and motivation
As part of our HSC AGN analysis, we undertake a pilot study for the search for rare massive black holes (MBHs) that are spatially offset from their host galaxy nuclei, in preparation for a larger-scale analysis with LSST. Recent cosmological simulations that do not artificially tie MBHs to halo centers have found that the majority of MBHs of mass $M_{\text{BH}}<10^{7}M_{\odot}$ have a $>15$kpc spatial offset from their host nucleus, with the fraction increasing for decreasing mass (Bellovary et al., 2019; Ricarte et al., 2021). Variability-selection has proven to be particularly effective at identifying low mass AGN in dwarf galaxies where traditional spectroscopic diagnostics are less effective (e.g. Baldassare et al., 2018, 2020; Ward et al., 2022; Burke et al., 2022), making it a promising approach for the identification of wandering MBHs. Forecasts for the number of $10^{4}M_{\odot}<M_{\text{BH}}<10^{6}M_{\odot}$ MBHs with detectable variability in LSST range from 1,500 to 21,000 depending on seed mass and wandering fraction (Burke et al., 2023).
There are a variety of predictions for the accretion luminosities of such wandering AGN (Pacucci et al., 2018; Di Matteo et al., 2023), but even if they are not observable from typical gas accretion, they may be observable as tidal disruption events if the wandering MBHs retain their nuclear star clusters (e.g Lin et al., 2018). The goal of this pilot study is to determine what spatial offsets will be detectable by applying multi-epoch scene modeling to LSST-like imaging data, using a sample of variable AGNs with well-determined spectroscopic redshifts and black hole masses (Aihara et al., 2022; Burke et al., 2024).
@@ 5.2 Light curves, host galaxy SEDs, and spatial offset distributions for variable AGN in HSC
In Figure 7, we show an HSC observation, the scarlet2 model for that observation derived from fitting the full multi-epoch data set, and the corresponding residual image for two COSMOS AGN: one in a compact, AGN-dominated galaxy – which is very typical for this sample – and one in an extended galaxy of comparable flux to the variable AGN. scarlet2 produces high quality models and light curves for all AGN in the HSC sample. Three example light curves from this procedure are shown in Figure 8.
@@
6 Technical Discussion
We tested the run times when undertaking modeling of a series of 100 multi-epoch ZTF images of size 120 by 120 pixels, containing one TDE, one host galaxy, and two background galaxies. Using an AMD EPYC 7763 CPU on the NERSC cluster Perlmutter [^2] https://docs.nersc.gov/systems/perlmutter/architecture/, producing the full scene model using the setup described in Section 4 took an average of 329.6 s over 10 runs. By comparison, by switching to the 40GB Nvidia A100 GPU, this time was reduced to 52.6 s, producing a 6.3 times speed-up. For cutouts that are not pre-aligned to the same wcs, the rendering operation is repeated throughout the fitting, increasing the CPU processing time from 5.5 minutes to 3.5 hours. It is therefore recommended that for large multi-epoch datasets, the imaging be pre-aligned if possible to reduce computation time. We also note that the GPU time for MCMC sampling over all flux, position and morphology parameters took 4.5 hours on average. Future work will implement a Hessian vector product approximation of the Fisher matrix enabling fast approximation of parameter uncertainties, allowing users to avoid full MCMC sampling unless necessary for detailed analysis of the posteriors.
We re-emphasize that, in most situations, difference imaging provides a faster method of transient detection and an optimal light curve for characterizing variability relative to a reference. However, a full scene model is advantageous when joint modeling of the host galaxy is required – such as for the measurement of transient-host offsets – and when the transient or AGN is already present in the reference imaging. A scene modeling approach provides a convenient way to jointly model imaging data taken with different telescopes at different times, enabling the production of light curves across multiple surveys without concern for reference image mismatch. The ability to simultaneously sample the posteriors of both transient and host galaxy parameters can be beneficial for detailed characterization of sources of interest.
In the HSC AGN analysis presented in Section 5, we chose to enforce a Sérsic profile for the galaxy morphologies when fitting. Adding this stronger constraint on galaxy morphology is only recommended for compact AGN host galaxies where there are strong degeneracies between flux from the variable AGN and the host. This constraint is not required for nuclear transients where pre-flare imaging is available, or for host galaxies where the extended galaxy emission beyond the central AGN is well-resolved in at least one image, enabling unambiguous measurement of the host galaxy SED. A carefully constructed prior for high redshift AGN host galaxies could also be used to avoid the need for a Sérsic profile constraint in this test case. We also note that for the AGN variability science cases described in Section 5.1, systematic flux errors introduced in the AGN light curve via the enforcement of a Sérsic profile do not inhibit the detection of the AGN variability itself, or analysis of the variability timescale.
@@ 7 Summary and conclusions
This method will be most powerful when high cadence time-domain ground-based imaging from surveys like ZTF and LSST is jointly modeled with higher resolution space-based imaging from the Hubble Space Telescope, Euclid (Racca et al., 2016), and the Roman Space Telescope (Akeson et al., 2019). The option for GPU acceleration means that image processing with scarlet2 can be run at scale. We have made the scarlet2 code available for the wider astrophysical community at https://github.com/pmelchior/scarlet2. Example jupyter notebooks for the methods described in each section are available at https://github.com/charlotteaward/TimeDomainScarlet2.
8 Acknowledgements
We thank the anonymous reviewers for their helpful feedback that was very beneficial to the paper. Charlotte Ward acknowledges support from the LSST Discovery Alliance under grant AWD1008640 (PI: Ward).